[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jun 17, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Jun 17 08:02:31 UTC 2017


HSU Shift Log for Sat Jun 17 2017 04:02:26 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: Several dead flying ants or termites in HSU spaces by  
bug lights; large roach in breakroom. Possible bug infestation developing  
due to recent heavy rains?
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170617/a31092f7/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list